Let Q1=x(1.1) ,Q2=x(1.2), Q3=x(1.3). Then Let Q= ln(3 +|Q1|+ 2|Q2|+ 3|Q3|), Then T= 5 sin2(100Q)

1) where x=x(t) solves x′′+x= tan(t), x(0) = 1, x′(0) = 2

2) where x=x(t) solves x′′−x=te^t, x(0) = 1, x′(0) = 2.

3) where x=x(t) solves x′′−x=t^2, x(0) = 1, x′(0) = 2

4) where x=x(t) solves x′′−2x′+x=(e^t/2t), x(1) = 1, x′(1) = 2

Please show all steps and thank you!
Answer #1

Y 4X= tant , x100=1, x1(0)=2. Auxiliary equation is given by mvf1zo * mati is complementary functions Cifo Geost +Casimzwe know particularm integral PPS - Jos hy did + 0,13 -- core faime trointestet one ſecond epoca.com 2- cost sina tant, dit Sivening given conditions, we get XCO)1 - 1- 9- 1x loglitol ef.921 X = 96-simt) of Ca cost - cost a (settentot sacz) + log Seed--tet, X0=1) 1109= £. bay as Auxiliary equation is given mrizo af met -Coff complementos y Aumestrom) o get of ce et Pute ok- $(-3+ -]+ se $ [taţ] 3 [ot] hence general soborone is given by Je Capit P.I [treating gex] & Xo Get ut eget teh hrt) using- qn Cove+t- „ - © & oto 7 24 21+ = 42 792 12 13 + 2 = 1 hence X(4)= 4yet Ete (19). cohreh in the requiered walue of xo3) x_x= tv, x(0)21, X1109=2. Auxiliary equation is given by mkizo 29 m afl is caps get thet P.R= Len DEI antoren 2 [1-Dw] lenUsing Condition tool. a la atca-2 946=3. - X(t) = Goto Creta et using X1 (6)=2 no es que2 - no 0+0=) 2455 * = 4 e) 5GF2 as 3-Here, * XI!_2014x-1 ), x(i)=1, X(X)=2. Auxiliary equation is givens wory mkem.+1=0 ay ( ) meln. Cipa Gregtet = 24. Dat langsusing X(1)= 1 1=16+C)e + o - G4C2 = te o X117)- (1409 )ett et (09) + b/1ggede xet ther.stat -7%!= (GtCot)et + etice + b/legedHena *(+)-betet + $t (1934at. which is the required value of a thence the result.